OpenStudy (toxicsugar22):

The center of a circle is the point C(2,4), and P(-1,6) is a point on the circle. Find the equation of the line tangent to the circle at the point P.

OpenStudy (anonymous):

Already answered: distance = sqrt (13) Answer: \[(x-2)^2 + (y-4)^2 = 13\]

OpenStudy (anonymous):

i think this question asks for tangent line yes? not equation of the circle. need calc.

OpenStudy (anonymous):

Lol. I fail. Answer will be up in a second. FORGIVE ME :D

OpenStudy (anonymous):

slope for CP = 6-4/ -1-2 = -2/3 slope for line (perpendicular to CP) = 3/2 y = mx +b m = 3/2 6 = -3/2 + b; b = 15/2 y = 3x/2 +15/2

OpenStudy (anonymous):

oops i fail too! i thought i would have to use implicit diff! F for me.

OpenStudy (anonymous):

Lol. Implicit diff would have been hard and it wouldn't have worked :D